# SnapMem: Snapshot-based 3D Scene Memory for Embodied Exploration and Reasoning

### v1

@@ 4.3 GOAT-Bench
On GOAT-Bench [17] (Table 3), we evaluate SnapMem’s effectiveness as a lifelong memory system that facilitates efficient exploration and reasoning. More implementation details are in Appendix 9.3.
@@
Baselines. Similar to the experiments in A-EQA, we compare SnapMem with Explore-EQA [29] and ConceptGraph [8] baselines. Due to implementation differences in Explore-EQA, we introduce an additional success criterion for this baseline: a subtask is considered successful if the target object is visible in the final observation. This extra criterion leverages ground truth grounding, thereby enhancing the baseline’s capability. To demonstrate the effectiveness of SnapMem’s lifelong memory, we include another baseline (SnapMem w/o memory) in which we clear the constructed scene graph after each subtask. We also directly include baselines implemented in GOAT-Bench. However, these baselines are simple RNN-based models trained via reinforcement learning, which causes their performance to lag behind the baselines we implemented.
Analysis. As shown in Table 3, SnapMem achieves the highest scores compared to previous methods in both accuracy and efficiency. Our major observations are as follows: 1) Even though GOAT-Bench is an object-based navigation benchmark, which is well-suited for ConceptGraph settings, SnapMem still outperforms ConceptGraph w/ frontier snapshots. This can be attributed to the snapshot-based representation, which captures more comprehensive information, making it easier to match with the diverse descriptions in GOAT-Bench. 2) When compared with the original SnapMem, the performance of SnapMem w/o memory declines for both GPT-4o and LLaVA-7B models, particularly in efficiency (SPL), indicating that SnapMem is beneficial as a memory system for lifelong learning. 3) Explore-EQA, which uses a traditional value map for each subtask to indicate regions of interest, also performs worse, as it lacks the mechanism to memorize information in explored regions. 4) In terms of efficiency, after each episode, SnapMem generates an average of 16.58 memory snapshots from 91.37 egocentric observations, and after prefiltering with $K=10$, only 4.66 memory snapshots are kept for each query.
@@
SnapMem w/o memory 40.6 14.6
SnapMem (Ours) 49.6 29.4
@@
SnapMem w/o memory 58.6 38.5
SnapMem (Ours) 69.1 48.9
@@ 5 Conclusion
We present SnapMem, a snapshot-based 3D scene memory that uses a set of informative snapshot images to cover the scene and store robust visual information. With the integration of the frontier-based exploration framework, SnapMem allows the agent to either leverage the memory of explored regions to solve tasks or explore the scene to expand its knowledge. With its incremental construction and efficient memory retrieval mechanism, SnapMem serves as an effective memory system for lifelong agents. Extensive experiments demonstrate the significant advantages of SnapMem over traditional scene representations.
@@ 6.1 Detailed Experiment Results
A-EQA. Table 4 presents a detailed breakdown of results on A-EQA across the seven OpenEQA question categories. As demonstrated in the table, SnapMem significantly outperforms ConceptGraph w/ Frontier Snapshots for questions requiring spatial reasoning, including spatial understanding and object localization where the relative positions of surroundings is needed to generate better answers. Such performance gain is attributed to Memory Snapshot, which visually stores both the foreground inter-object spatial relationships and background room-level spatial cues. In contrast, ConceptGraph relies solely on object-centric representations, limiting its ability to capture broader spatial context. For other question categories focus on identifying object-specific variables, i.e., object recognition, attribute recognition, object state recognition, or heavily rely on external knowledge embedded within VLMs, i.e., world knowledge, SnapMem also showcases comparable performance as it ensures the capture of all informative objects and effectively utilizes the capability of VLMs.
@@
Compared with Explore-EQA, SnapMem generally exhibits higher LLM-Match scores in object-related question categories as we explicitly represents major objects within the scene by Memory Snapshots, enabling the agent to concentrate on relevant elements that may contribute to the final answer. On the other hand, Explore-EQA has consistently lower SPL due to its inefficient semantic-map-based exploration mechanism where explicit visual information of frontiers is encoded into an implicit semantic map. SnapMem addresses this limitation by visually capturing glimpses of unexplored areas with Frontier Snapshots and integrating them with Memory Snapshots in the decision-making phase, which provides a more intuitive and holistic view, enabling it to make more informed and effective choices during exploration.
@@
GOAT-Bench. Table 5 presents a detailed breakdown of results on GOAT-Bench across the three question modalities. Comparing SnapMem with CG w/ Frontier Snapshots, we observe that SnapMem significantly outperforms CG in the Object Category and Language modalities. This improvement is attributed to SnapMem’s memory snapshots, which provide explicit spatial relationships among objects and their surroundings, enabling the agent to locate targets more effectively. The detailed spatial context captured in the snapshots enhances the agent’s ability to interpret instructions that rely on spatial cues. In contrast, SnapMem’s SPL in the Image modality is slightly lower than CG’s, despite a similar Success Rate. This decrease is likely due to current vision-language models (VLMs) struggling to relate images of complex scenes taken from different angles. When the memory snapshots and the image prompts depict the same region from different perspectives, the VLM may become distracted, leading to less efficient navigation paths. This may highlight a limitation in current VLMs’ ability to match images across varying viewpoints in complex environments.
@@
SnapMem consistently outperforms both methods across all modalities in terms of Success Rate and SPL scores. By enabling the agent to recall previously observed regions and objects, memory significantly enhances the effectiveness and efficiency of exploration and reasoning. These results highlight memory’s essential role in lifelong object navigation tasks.
@@
SnapMem w/o memory 55.6 16.0 33.3 15.5 31.8 12.2 40.6 14.6
SnapMem (Ours) 62.6 33.3 49.5 31.7 35.2 22.7 49.6 29.4
@@
SnapMem w/o memory 69.9 45.4 50.35 30.1 54.4 39.5 58.6 38.5
SnapMem (Ours) 79.2 55.8 61.9 46.0 65.2 44.2 69.1 48.9
@@ 7 Failure Case Analysis
In Figure 5 to 10, we analyze and categorize the types of questions where SnapMem performs poorly in A-EQA. For each example question, we provide the ground truth answer, SnapMem’s predicted answer, and the memory snapshot selected for answering the question. Each memory snapshot includes object detection annotations to better visualize which objects are detected and incorporated into the scene graph. These annotations can be zoomed in for a clearer view. Note that these annotations are not part of the actual input provided to the VLM.
@@
  • Limitations of the Object Detection Model. SnapMem relies on an object detector to identify and add new objects to the scene graph. However, the object detector can sometimes produce incorrect labels, as shown in Figure 8. In most cases, the prefiltering process successfully filters out these incorrect labels, as they are often highly irrelevant. Additionally, the VLM is generally capable of recognizing and ignoring such errors. However, certain situations, as in Figure 9, illustrate cases where the detector mislabels objects—such as detecting a TV as a fan or a cloth rack as a ladder. These misclassifications can confuse the VLM, especially when the incorrect labels closely align with the expected answer. In addition to incorrect detections, the target objects for answering the questions are detected at all. As shown in Figure 9, if the car or the monitor had been detected in the relevant snapshots, the question could have been answered correctly. However, since these objects were not detected and included in memory, the VLM could not select an appropriate memory snapshot, and the agent eventually exceeded the step limit. Interestingly, in some cases, as shown in Figure 10, even when the target objects are not detected, they remain visible in other memory snapshots that pass the prefiltering process. In these cases, the VLM is still able to answer the question successfully. This demonstrates that SnapMem is more robust than traditional 3D scene graph approaches.
@@ 8 Details of Frontier-based Exploration Framework
Our frontier-based exploration framework is based on the framework in Explore-EQA [29]. We enhance its robustness and adapt it to our snapshot-based representation framework. A 3D grid-based occupancy map $M$, representing the length, width and height of the entire room, is used to record the occupancy, with each voxel having a side length of 0.1 meters. During exploration, each depth observation, together with its corresponding observation pose, is used to map unoccupied spaces onto the initially fully occupied $M$. The navigable region is then defined as the layer of unoccupied voxels at the height of 0.4 meters above the ground where the agent moves. Within this navigable region, the area within 1.7 meters of the agent’s trajectory is defined as the explored region, while the remainder is designated as the unexplored region, as illustrated in Figure 11.
@@ 9.1 Implementation Details for A-EQA
As explained in detail in Section 3.2, we integrate SnapMem into the frontier-based exploration framework. The VLM directly returns an answer after identifying visual clues from certain memory snapshots. We set the number of egocentric observations at each step $N=3$, the maximum distance for objects to be included in the scene graph $max\_dist=3.5$, and the number of prefiltered classes $K=10$.
@@ 9.2 Implementation Details for EM-EQA
To adapt SnapMem to the EM-EQA benchmark, we first construct SnapMem for each scene using the given RGB-D observations and corresponding camera poses. For each question, we then apply prefiltering to the memory snapshots using different K values (1, 2, 3, 5, 10), and utilize the resulting filtered snapshots as prompts for GPT-4o to generate the answers.
@@ 10 Details of the Active Exploration
Figure 12: Overview of the frontier-based exploration pipeline with SnapMem on embodied question-answering task.
@@ 12 Ablation Study
In Figure 15, we present the evaluation metrics for different choices of $K$ on both A-EQA and GOAT-Bench. In addition to the metrics introduced in the experiment sections, we include the average ratio of the number of remaining memory snapshots after prefiltering to the total number of memory snapshots as a measure of the effectiveness and intensity of prefiltering. The results on both benchmarks align with our intuition: allowing more prefiltered classes leads to better performance. Moreover, even when $K=10$, on average only 3.26 and 4.66 memory snapshots are left after prefiltering for A-EQA and GOAT-Bench respectively, accounting for 29.8% and 28.1% of the total memory snapshots, and 8.2% and 5.1% of the total frame candidates. These statistics demonstrate the effectiveness of prefiltering as a memory retrieval mechanism, as well as SnapMem’s compactness as a scene representation. Furthermore, we observe that the overall performance does not drop significantly when $K$ is small, highlighting the robustness of our framework.
@@ 13 Other Related Works
Our proposed SnapMem first differs from these topological mapping methods in its focus and design. SnapMem focuses on capturing all salient objects in the scene by a minimum number of memory snapshots. Each memory snapshot is designed to capture the visual features of a cluster of objects in the nearby region, along with their spatial relationships and surrounding environment. Objects are uniquely assigned to one memory snapshot, making the representation informative, comprehensive, and compact—key qualities for leveraging vision-language models (VLMs) with limited context length to interpret and reason over visual data. In contrast, the images in topological map in [18] are primarily designed to represent landmarks for navigation, without attempting to capture all informative aspects of the scene or focusing on visually representing all objects in a 3D environment. The images in the representation in [6], though focused on object segments, are not compact, containing redundancy between consecutive frames, and the images still serve as navigation landmarks rather than visual representations of inter-object relationships.
In addition, SnapMem introduces the concept of frontier snapshots to explicitly model unexplored regions, allowing agents to make informed decisions about where to explore next to expand knowledge—an active exploration capability not addressed in previous 2D methods. Moreover, the structure of SnapMem enables the memory retrieval mechanism of Prefiltering, which manages the memory scalability and efficiency over extended operations, supporting lifelong learning that is absent in the aforementioned works. Lastly, like other 3D scene graphs, SnapMem stores the 3D information of the objects and snapshots. Based on this information, as in the practice of ConceptFusion [15], a set of spatial relationship comparators can be called by LLMs as queries, e.g., querying the distance between A and B by calling “howFar(A, B)”. This information is cannot be stored in those 2D representations.

### v2

@@ 4.3 GOAT-Bench
On GOAT-Bench [17] (Table 3), we evaluate 3D-Mem’s effectiveness as a lifelong memory system that facilitates efficient exploration and reasoning. More implementation details are in Appendix 9.3.
@@
Baselines. Similar to the experiments in A-EQA, we compare 3D-Mem with Explore-EQA [29] and ConceptGraph [8] baselines. Due to implementation differences in Explore-EQA, we introduce an additional success criterion for this baseline: a subtask is considered successful if the target object is visible in the final observation. This extra criterion leverages ground truth grounding, thereby enhancing the baseline’s capability. To demonstrate the effectiveness of 3D-Mem’s lifelong memory, we include another baseline (3D-Mem w/o memory) in which we clear the constructed scene graph after each subtask. We also directly include baselines implemented in GOAT-Bench. However, these baselines are simple RNN-based models trained via reinforcement learning, which causes their performance to lag behind the baselines we implemented.
Analysis. As shown in Table 3, 3D-Mem achieves the highest scores compared to previous methods in both accuracy and efficiency. Our major observations are as follows: 1) Even though GOAT-Bench is an object-based navigation benchmark, which is well-suited for ConceptGraph settings, 3D-Mem still outperforms ConceptGraph w/ frontier snapshots. This can be attributed 3D-Mem’s multi-view images representation, which captures more comprehensive information, making it easier to match with the diverse descriptions in GOAT-Bench. 2) When compared with the original 3D-Mem, the performance of 3D-Mem w/o memory declines for both GPT-4o and LLaVA-7B models, particularly in efficiency (SPL), indicating that 3D-Mem is beneficial as a memory system for lifelong learning. 3) Explore-EQA, which uses a traditional value map for each subtask to indicate regions of interest, also performs worse, as it lacks the mechanism to memorize information in explored regions. 4) In terms of efficiency, after each episode, 3D-Mem generates an average of 16.58 memory snapshots from 91.37 egocentric observations, and after prefiltering with $K=10$, only 4.66 memory snapshots are kept for each query.
@@
3D-Mem w/o memory 40.6 14.6
3D-Mem (Ours) 49.6 29.4
@@
3D-Mem w/o memory 58.6 38.5
3D-Mem (Ours) 69.1 48.9
@@ 5 Conclusion
We present 3D-Mem, a 3D scene memory framework that uses a set of informative multi-view snapshot images to cover the scene and store robust visual information. With the integration of the frontier-based exploration framework, 3D-Mem allows the agent to either leverage the memory of explored regions to solve tasks or explore the scene to expand its knowledge. With its incremental construction and efficient memory retrieval mechanism, 3D-Mem serves as an effective memory system for lifelong agents. Extensive experiments demonstrate the significant advantages of 3D-Mem over traditional scene representations.
@@ 6.1 Detailed Experiment Results
A-EQA. Table 4 presents a detailed breakdown of results on A-EQA across the seven OpenEQA question categories. As demonstrated in the table, 3D-Mem significantly outperforms ConceptGraph w/ Frontier Snapshots for questions requiring spatial reasoning, including spatial understanding and object localization where the relative positions of surroundings is needed to generate better answers. Such performance gain is attributed to Memory Snapshot, which visually stores both the foreground inter-object spatial relationships and background room-level spatial cues. In contrast, ConceptGraph relies solely on object-centric representations, limiting its ability to capture broader spatial context. For other question categories focus on identifying object-specific variables, i.e., object recognition, attribute recognition, object state recognition, or heavily rely on external knowledge embedded within VLMs, i.e., world knowledge, 3D-Mem also showcases comparable performance as it ensures the capture of all informative objects and effectively utilizes the capability of VLMs.
@@
Compared with Explore-EQA, 3D-Mem generally exhibits higher LLM-Match scores in object-related question categories as we explicitly represents major objects within the scene by Memory Snapshots, enabling the agent to concentrate on relevant elements that may contribute to the final answer. On the other hand, Explore-EQA has consistently lower SPL due to its inefficient semantic-map-based exploration mechanism where explicit visual information of frontiers is encoded into an implicit semantic map. 3D-Mem addresses this limitation by visually capturing glimpses of unexplored areas with Frontier Snapshots and integrating them with Memory Snapshots in the decision-making phase, which provides a more intuitive and holistic view, enabling it to make more informed and effective choices during exploration.
@@
GOAT-Bench. Table 5 presents a detailed breakdown of results on GOAT-Bench across the three question modalities. Comparing 3D-Mem with CG w/ Frontier Snapshots, we observe that 3D-Mem significantly outperforms CG in the Object Category and Language modalities. This improvement is attributed to 3D-Mem’s memory snapshots, which provide explicit spatial relationships among objects and their surroundings, enabling the agent to locate targets more effectively. The detailed spatial context captured in the snapshots enhances the agent’s ability to interpret instructions that rely on spatial cues. In contrast, 3D-Mem’s SPL in the Image modality is slightly lower than CG’s, despite a similar Success Rate. This decrease is likely due to current vision-language models (VLMs) struggling to relate images of complex scenes taken from different angles. When the memory snapshots and the image prompts depict the same region from different perspectives, the VLM may become distracted, leading to less efficient navigation paths. This may highlight a limitation in current VLMs’ ability to match images across varying viewpoints in complex environments.
@@
3D-Mem consistently outperforms both methods across all modalities in terms of Success Rate and SPL scores. By enabling the agent to recall previously observed regions and objects, memory significantly enhances the effectiveness and efficiency of exploration and reasoning. These results highlight memory’s essential role in lifelong object navigation tasks.
@@
3D-Mem w/o memory 55.6 16.0 33.3 15.5 31.8 12.2 40.6 14.6
3D-Mem (Ours) 62.6 33.3 49.5 31.7 35.2 22.7 49.6 29.4
@@
3D-Mem w/o memory 69.9 45.4 50.35 30.1 54.4 39.5 58.6 38.5
3D-Mem (Ours) 79.2 55.8 61.9 46.0 65.2 44.2 69.1 48.9
@@ 7 Failure Case Analysis
In Figure 5 to 10, we analyze and categorize the types of questions where 3D-Mem performs poorly in A-EQA. For each example question, we provide the ground truth answer, 3D-Mem’s predicted answer, and the memory snapshot selected for answering the question. Each memory snapshot includes object detection annotations to better visualize which objects are detected and incorporated into the scene graph. These annotations can be zoomed in for a clearer view. Note that these annotations are not part of the actual input provided to the VLM.
@@
  • Limitations of the Object Detection Model. 3D-Mem relies on an object detector to identify and add new objects to the scene graph. However, the object detector can sometimes produce incorrect labels, as shown in Figure 8. In most cases, the prefiltering process successfully filters out these incorrect labels, as they are often highly irrelevant. Additionally, the VLM is generally capable of recognizing and ignoring such errors. However, certain situations, as in Figure 9, illustrate cases where the detector mislabels objects—such as detecting a TV as a fan or a cloth rack as a ladder. These misclassifications can confuse the VLM, especially when the incorrect labels closely align with the expected answer. In addition to incorrect detections, the target objects for answering the questions are detected at all. As shown in Figure 9, if the car or the monitor had been detected in the relevant snapshots, the question could have been answered correctly. However, since these objects were not detected and included in memory, the VLM could not select an appropriate memory snapshot, and the agent eventually exceeded the step limit. Interestingly, in some cases, as shown in Figure 10, even when the target objects are not detected, they remain visible in other memory snapshots that pass the prefiltering process. In these cases, the VLM is still able to answer the question successfully. This demonstrates that 3D-Mem is more robust than traditional 3D scene graph approaches.
@@ 8 Details of Frontier-based Exploration Framework
Our frontier-based exploration framework is based on the framework in Explore-EQA [29]. We enhance its robustness and adapt it to our multi-view images representation framework. A 3D grid-based occupancy map $M$, representing the length, width and height of the entire room, is used to record the occupancy, with each voxel having a side length of 0.1 meters. During exploration, each depth observation, together with its corresponding observation pose, is used to map unoccupied spaces onto the initially fully occupied $M$. The navigable region is then defined as the layer of unoccupied voxels at the height of 0.4 meters above the ground where the agent moves. Within this navigable region, the area within 1.7 meters of the agent’s trajectory is defined as the explored region, while the remainder is designated as the unexplored region, as illustrated in Figure 11.
@@ 9.1 Implementation Details for A-EQA
As explained in detail in Section 3.2, we integrate 3D-Mem into the frontier-based exploration framework. The VLM directly returns an answer after identifying visual clues from certain memory snapshots. We set the number of egocentric observations at each step $N=3$, the maximum distance for objects to be included in the scene graph $max\_dist=3.5$, and the number of prefiltered classes $K=10$.
@@ 9.2 Implementation Details for EM-EQA
To adapt 3D-Mem to the EM-EQA benchmark, we first construct 3D-Mem for each scene using the given RGB-D observations and corresponding camera poses. For each question, we then apply prefiltering to the memory snapshots using different K values (1, 2, 3, 5, 10), and utilize the resulting filtered snapshots as prompts for GPT-4o to generate the answers.
@@ 10 Details of the Active Exploration
Figure 12: Overview of the frontier-based exploration pipeline with 3D-Mem on embodied question-answering task.
@@ 12 Ablation Study
In Figure 15, we present the evaluation metrics for different choices of $K$ on both A-EQA and GOAT-Bench. In addition to the metrics introduced in the experiment sections, we include the average ratio of the number of remaining memory snapshots after prefiltering to the total number of memory snapshots as a measure of the effectiveness and intensity of prefiltering. The results on both benchmarks align with our intuition: allowing more prefiltered classes leads to better performance. Moreover, even when $K=10$, on average only 3.26 and 4.66 memory snapshots are left after prefiltering for A-EQA and GOAT-Bench respectively, accounting for 29.8% and 28.1% of the total memory snapshots, and 8.2% and 5.1% of the total frame candidates. These statistics demonstrate the effectiveness of prefiltering as a memory retrieval mechanism, as well as 3D-Mem’s compactness as a scene representation. Furthermore, we observe that the overall performance does not drop significantly when $K$ is small, highlighting the robustness of our framework.
@@ 13 Other Related Works
Our proposed 3D-Mem first differs from these topological mapping methods in its focus and design. 3D-Mem focuses on capturing all salient objects in the scene by a minimum number of memory snapshots. Each memory snapshot is designed to capture the visual features of a cluster of objects in the nearby region, along with their spatial relationships and surrounding environment. Objects are uniquely assigned to one memory snapshot, making the representation informative, comprehensive, and compact—key qualities for leveraging vision-language models (VLMs) with limited context length to interpret and reason over visual data. In contrast, the images in topological map in [18] are primarily designed to represent landmarks for navigation, without attempting to capture all informative aspects of the scene or focusing on visually representing all objects in a 3D environment. The images in the representation in [6], though focused on object segments, are not compact, containing redundancy between consecutive frames, and the images still serve as navigation landmarks rather than visual representations of inter-object relationships.
In addition, 3D-Mem introduces the concept of frontier snapshots to explicitly model unexplored regions, allowing agents to make informed decisions about where to explore next to expand knowledge—an active exploration capability not addressed in previous 2D methods. Moreover, the structure of 3D-Mem enables the memory retrieval mechanism of Prefiltering, which manages the memory scalability and efficiency over extended operations, supporting lifelong learning that is absent in the aforementioned works. Lastly, like other 3D scene graphs, 3D-Mem stores the 3D information of the objects and snapshots. Based on this information, as in the practice of ConceptFusion [15], a set of spatial relationship comparators can be called by LLMs as queries, e.g., querying the distance between A and B by calling “howFar(A, B)”. This information is cannot be stored in those 2D representations.